OpenStudy (anonymous):

Can someone please help me with this question Locate the foci of the ellipse. Show your work. x^2/36+y^2/11=1

OpenStudy (alekos):

General form is x^2/a^2 + y^2/b^2=1 Where a^2=36 and b^2 =11 C^2 = a^2 - b^2 Where c= distance of focal points from the centre

OpenStudy (anonymous):

so what do i do @alekos

OpenStudy (anonymous):

just plug and play

OpenStudy (alekos):

Yes. Plug and play

OpenStudy (anonymous):

@dan815 hey can you tell me what do do next after this

OpenStudy (anonymous):

@Arguedausti793 if you could help me it would be much appreciated

OpenStudy (anonymous):

I'm sorry! I don't remember how to do this..

OpenStudy (anonymous):

ok

myininaya (myininaya):

@quick24 what part are you having problems with ? \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1 \text{ has center } (0,0)\] we want to find c and we can do that by Pythagorean theorem \[b^2+c^2=a^2 \\ \text{ or solving for } c^2 \text{ gives } \\ c^2=a^2-b^2 \\ \text{ then solving for c} \\ c=\sqrt{a^2-b^2}\]